OpenStudy (anonymous):

What is the new volume of a rectangular prism that is originally twelve cubic meters and is then resized using a scale factor of 3 on each side length? four cubic meters thirty-six cubic meters one hundred eight cubic meters three hundred twenty-four cubic meters

OpenStudy (jdoe0001):

 \(\bf 12=x\cdot x\cdot x=x^3\qquad {\color{red}{ 3}}x\cdot {\color{red}{ 3}}x\cdot {\color{red}{ 3}}x\implies {\color{red}{ 3^3}}x^3\\ \quad \\ x^3=12\qquad \qquad {\color{red}{ 3^3}}x^3={\color{red}{ 3^3}}\cdot 12\)